OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f ° g)(-5).

OpenStudy (anonymous):

can you find g(-5) = ????

OpenStudy (anonymous):

after you find g(-5), put that answer into f...

OpenStudy (anonymous):

I'm a little lost, so pretty much just do this, -2*(-5) -7 and -4*(-5) + 3? Is that what your saying?

OpenStudy (anonymous):

yes... but i only wanted g(-5)... so g(-5) = -4*(-5) + 3 = ???? what's the answer here?

OpenStudy (anonymous):

23

OpenStudy (anonymous):

good... now that you found g(-5)=23, that 23 goes into f. so you need to find f(23) = ???

OpenStudy (anonymous):

It would be -53 wouldnt it?

OpenStudy (anonymous):

that is the definition of (f ° g)(-5) = f(g(-5))

OpenStudy (anonymous):

yes.... :) f(23) = -2*(23) - 7 = -46 -7 = -53

OpenStudy (anonymous):

good work... :)

OpenStudy (anonymous):

Thanks for your help.! =)

OpenStudy (anonymous):

yw...:)